[Brahms-l] Shift report 20031231 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Wed Dec 31 2003 - 08:04:41 EST
------------------------------------------------------------
Date: Wed Dec 31 08:04:41 2003, Shift: 00:00-08:00
Supervisor: Catalin Ristea, with: 
Log book#15 pages 19 - 19

Shift Summary:
------------------------------------------------------------
1. Had beam from 2am to 4,30am. Some problems with the blue storage cavity.
2. Run 9284 - D1=D2=1/4A, D3=D4=0, D5=350A - 90k(tr1).
3. Some problems with the magnets (changing pol on D1)
   The DCs are up and running.
   RICH is off at this moment.

I take this opportunity to wish you all Brahms-ers a very good new year!
Thank you!

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Dec 31 08:05:32 2003

This archive was generated by hypermail 2.1.8 : Wed Dec 31 2003 - 08:05:45 EST